# SSupplemental References

### v2

@@ 2.2.1 Ranking method
To begin the AHP process, we constructed a pairwise comparison matrix (A), where each element $a_{ij}$ represented the importance of criterion $i$ relative to criterion $j$. For our example, the matrix was as follows:
$A=\begin{bmatrix}1&2&4&5&7&9\\ 0.5&1&3&4&5&7\\ 0.25&0.33&1&2&4&5\\ 0.2&0.25&0.5&1&3&5\\ 0.14&0.2&0.25&0.33&1&4\\ 0.11&0.14&0.2&0.2&0.25&1\end{bmatrix}$ (1)
@@
We then checked the consistency of the pairwise comparison matrix by calculating the Consistency Index (CI) and the Consistency Ratio (CR) (Saaty, 1980). The CI was calculated using the formula:
@@
In our case, with a consistency ratio (CR) of 0.06, the matrix was considered consistent (as CR < 0.1). This consistency indicated that our pairwise comparison matrix was reliable for determining the weights. To create the vulnerability map of the Ganga buffer region, if
  • $W$ is the 1x6 row vector of AHP weights:
    $W=\begin{bmatrix}w_{1}&w_{2}&w_{3}&w_{4}&w_{5}&w_{6}\end{bmatrix}$ (6)
We calculated the predicted vulnerability for each pixel by multiplying each factor’s parameter value by its corresponding AHP weight i.e. the dot product of the vector W with the transpose of the vector X(p):
$V(p)=W\cdot X(p)^{T}$ (8)
We used Google Earth Engine to multiply the AHP weights with the bucketed values for each layer at each pixel and aggregated them to get our final vulnerability scores for the entire Ganga buffer. More information about the Remote Sensing datasets we used on Google Earth Engine can be found in Supplemental Table S2.
2.3 Diagnostic Analyses
2.3.1 Nested AHP
We performed the AHP analysis under the assumption that while each parameter contributes to the vulnerability differently, the contribution of each parameter to vulnerability was linear in nature i.e. areas where the value of a parameter was 5 contributed to the vulnerability 5 times more than areas where the value was 1. This is an overly simplified approach; hence, we created a pairwise matrix for each factor, using the bucketed values for the parameters, and performed AHP on each parameter with itself, to create the Nested AHP approach.
@@ 2.3.1 Nested AHP
Using a pairwise comparison matrix, we assessed the relationship between individual parameter values. Table 2 shows the pairwise comparison of Population Density and the weight each category has, e.g. PD of >4036 people per km2 has a 9 times greater effect on vulnerability than when PD was <865 people per km2. The pairwise matrices for the other factors alongside the explanations of the comparison indices were provided in the Supplemental Information section.
Table 2: Nested AHP: Pairwise Comparison for Individual Parameter Values for Population Density
Population Density / km2 >4036 2680-4036 1600-2680 865-1600 <865
>4036 1 4 6 8 9
2680-4036 0.25 1 2 4 5
1600-2680 0.17 0.5 1 2 3
865-1600 0.13 0.25 0.5 1 1
<865 0.11 0.2 0.33 1 1
Total 1.65 5.95 9.83 16 19
@@
Running AHP on this pairwise matrix for Population Density gave the subweights for classes 1 - 5 (Table 3). Similarly, once we ran this for the other factors; LULC, Slope, Precipitation, Drainage Density, and Temperature, we got the priority vector of subweights for each factor.
Table 3: Nested AHP: Priority Vector of Subweights for each factor
@@
  • $S$ is the 5x6 matrix of subweights, where each column corresponds to a factor and each row corresponds to the subweights for factor classes 1, 2, 3, 4, 5 (as can be seen in Table 1).
@@
2.3.2 Fuzzy AHP
Fuzzy Analytic Hierarchy Process (Fuzzy AHP), as introduced by Saaty (1982) is an extension of the classic AHP that incorporates the fuzziness (probabilistic alterations/error) associated with human judgment. By using fuzzy numbers in the rankings, Fuzzy AHP accommodates these uncertainties and thus provides more reliable results.
To understand the robustness of our AHP factor rank order results, we utilized fuzzy AHP to see if there was an inconsistency within the results, i.e. if there was a rank reversal between the parameters when we incorporate fuzziness to the values of the AHP pairwise comparison matrix.
@@ 2.3.2 Fuzzy AHP
In Fuzzy AHP, the decision maker generally recreates the pairwise matrix by choosing the $l,m,u$ for each element, but to understand the outlying cases for Fuzzy AHP, and to incorporate the breadth of decision maker’s range, we ran Fuzzy AHP on 100,000 Monte Carlo simulations at 95% fuzziness. For each run, we made a unique Fuzzy Pairwise Comparison Matrix, where if the comparison index between two factors was 2, the triangular fuzzy range would be (0.05*2, 2, 2*1.95) = (0.1, 2, 3.9), and we chose a random value between 0.1 and 2, and 2 and 3.9 to design the final TFN [random.uniform(0.1,2), 2, random.uniform(2, 3.9)]. Running 100,000 simulations took into account most of the likely Fuzzy Pairwise Comparison Matrices that might be generated.
@@
2.3.3 ANP
One of the main limitations of AHP is its inability to adequately capture the interdependencies among criteria and alternatives. To address this issue, Saaty (2006) developed ANP, which extends the AHP methodology by accommodating feedback loops and interdependencies among criteria and alternatives. Unlike AHP, which assumes a hierarchical structure, ANP allows for the representation of complex relationships and feedback loops in decision networks. The method utilizes a supermatrix to integrate both local and global influences, providing a more realistic representation of environments. ANP is widely used to represent interdependent relationships, making it suitable for decision problems with interconnected elements. ANP introduces the concept of contextual weights, enabling decision makers to adjust the importance of criteria and alternatives based on the context of the decision problem. This flexibility enhances the adaptability of ANP to various situations.
In a complex system such as the geophysical environment, it is expected that AHP factors will be correlated to one another. While AHP assumes strict independence between criteria at different levels of the hierarchy, ANP allows for both dependence and independence relationships (Table 4) as described previously by Poh and Liang (2017); this flexibility better captures the intricacies of decision networks across natural systems.
Table 4: Vulnerability ANP Supermatrixa
Goal Criteria Subcriteria Alternative outcomes
Goal 0 0 0 0
Criteria (AHP factors) W21 W22 0 0
Subcriteria (vulnerability classes) 0 W32 W33 0
Alternative outcomes 0 0 W34 I
The ANP supermatrix component matrices as follows:
  • Relative importance of criteria (AHP factors), $W_{21}$;
  • Inner dependence (As defined by Poh and Liang (2017)) matrix of criteria, $W_{22}$;
  • Inner dependence matrix of subcriteria, $W_{33}$;
  • and Relative importance of subcriteria w.r.t alternatives, $W_{34}$.
$W_{21}$ $=\begin{bmatrix}0.408\\ 0.268\\ 0.138\\ 0.1\\ 0.058\\ 0.028\\ \end{bmatrix}\quad$ $W_{22}$ $=\begin{bmatrix}0.464&0.233&0.14&0.277&0&0.186\\ 0.271&0.408&0.12&0.098&0&0.077\\ 0.105&0.126&0.662&0&0&0\\ 0.085&0.079&0&0.498&0&0\\ 0.076&0.079&0.078&0.126&0&0\\ 0&0.075&0&0&0&0.737\\ \end{bmatrix}$
$W_{32}$ $=\begin{bmatrix}0.572&0.505&0.501&0.642&0.388&0.557\\ 0.207&0.186&0.172&0.157&0.233&0.165\\ 0.114&0.144&0.124&0.071&0.276&0.129\\ 0.057&0.11&0.102&0.065&0.07&0.099\\ 0.05&0.055&0.102&0.065&0.033&0.05\\ \end{bmatrix}\quad$ $W_{33}=O_{5x5}$
$W_{34}$ $=\begin{bmatrix}0.56&0.61&0.571&0.483&0.409\\ 0.187&0.203&0.229&0.276&0.273\\ 0.112&0.102&0.114&0.138&0.182\\ 0.08&0.051&0.057&0.069&0.091\\ 0.062&0.034&0.029&0.034&0.045\\ \end{bmatrix}\qquad$ $I=I_{5x5}$
The calculation of each component matrix is available in the “ANP supermatrix and inner dependence pairwise comparisons” section in the Supplemental Information.
We generated the ANP supermatrix with the goal to assign pollution vulnerability scores (PVS). Here the AHP Factors, or Criteria were - Population density, LULC, rainfall, drainage density, slope and temperature. The Alternatives were pertaining to the PVS, i.e., extremely high, high, moderate, low, extremely low. The possible relationship arcs were dependencies, self-loops, and feedbacks.
We used the Python library pyanp created by Adams et al. (2020) to run the ANP Algorithm on the Supermatrix.
2.3.4 1-N AHP
Another shortcoming of AHP is that, from inside the process, there is no way of knowing for certain if there are any important factors missing from the decision problem parameters. For AHP, our model was considering six factors - Population density, LULC, rainfall, drainage density, slope and temperature. If a fundamental pollution vulnerability factor was missing, AHP would not capture it.
To overcome this, we devised an approach named 1-N AHP where we aimed to validate whether the unknown and unconsidered factors are important for our analysis of vulnerability or not.
@@ 2.3.3 ANP
We first took our original output of AHP, i.e. the weights of the known factors. Here, we assumed that our known factors would at least represent 67% of the importance while classifying the vulnerability variable. This is a fair assumption since through our preliminary analysis we can conclude that our factors were robust enough to represent at least 67% of the variability of the vulnerability variable.
@@
To evaluate the unknown variability captured between 1-N (where N lies between 0.67 and 1), we split it between acute and chronic. We assumed that since acute factors have a significant impact on vulnerability, 90% of the 1-N factor would be attributed to the impact of acute factors to vulnerability, while the leftover 10% would be attributed to the impact of chronic factors to vulnerability. For instance, in the worst case scenario if 1-N is 0.33 (i.e. the weightage of the unknown factors towards vulnerability is 33%), the 0.33 will be split into 0.297 for acute and 0.033 for chronic. The original six factors, with their weights 0.408, 0.268, 0.138, 0.100, 0.058, 0.028 will be scaled down based on N. If N is 0.67, we multiply the old weights by 0.67 so that they sum up to n, and we can accommodate the new acute and chronic weights. Here, 0.408*0.67 gives 0.27336, which would be lower than 0.297 for acute factors, which means the acute factors would have a higher impact on vulnerability but with a lower probability.
@@
2.3.5 Fuzzy 1-N AHP
We have described two approaches to deal with the complex nature of decision making methods for environmental problems. While Fuzzy AHP took into account the potential error in the decision maker’s choice for the values of the pairwise comparison matrix, 1-N AHP considered unknown factors which we may have missed while considering the vulnerability of a river. Since both approaches were mutually exclusive, we performed one final test to check the robustness of our results by combining these two approaches.
We performed 1-N AHP followed by a fuzzy analysis on the values. First, we used our eight weights (six original weights, plus a weight each for aggregated acute and chronic factors) that we derived from our analysis in 1-N AHP, and applied linear algebra to create our new matrix (i.e. because the AHP vector output is the eigenvector, and the eigenvalue is used to calculate the confidence, we used $A=PDP^{-1}$ to get back our original matrix $A$, where $P$ is a matrix of eigenvectors and $D$ is a diagonal matrix of eigenvalues). We then ran Fuzzy AHP on this new 8x8x3 matrix for 100,000 simulations to get the 8x1 size priority vectors. From these vectors, we randomly excluded the acute factors for 97.5% of the cases, and in 75% of the 100,000 cases, the chronic factors were randomly excluded, before checking the rank reversals. This was to account for the probabilities of the presence (or absence) of acute and chronic factors respectively while checking for rank reversals.
3 Results
3.1 River Ganga vulnerability based on AHP
@@
Fig 3. These figures show the AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
AHP vulnerability scores ranged from a minimum of 1.001 to a maximum of 4.470 with 83.7% of the area being categorized as having extremely low or low vulnerability and 3.5% of the area is highly or extremely high vulnerability (Fig 3b). Tier 2 cities (i.e., cities with a population between 50,000 and 99,999 people) on the banks of river Ganga such as Varanasi, Prayagraj/Allahabad, and Kanpur were clearly highlighted with high and extremely high AHP vulnerability scores (Fig 3a). Indeed, urban settings that had population densities $\geq$ 1,100 people km2 typically had vulnerability scores $\geq$ 2.736. Alternatively, the areas with vulnerability scores $\leq$ 1.694 were characterized by forest or cropland regions with low human population densities. The annual rainfall in the high elevation regions of the river Ganga generally increased the vulnerability of those regions.
3.2 Nested AHP and ANP
The distribution of vulnerability scores produced with both Nested AHP and ANP were markedly different from the distribution produced with AHP (Fig 4 and Fig 5). Specifically, Nested AHP resulted in 96.7% of the area scored as extremely low and low vulnerability, which was skewed with a right-tail range that extended 1.8-fold further than AHP. Similarly, ANP exhibited a right-tailed distribution with low vulnerability scores. Despite that, the Tier 2 cities had high vulnerability scores for both ANP and Nested AHP, except for Kanpur, which saw a major reduction in the vulnerability scores. These differences were due to the nonlinear response of factor weights to vulnerability rankings in both methods (Fig 6). Thus, these methods resulted in overall focused, yet dampened, vulnerability-score maps compared to AHP. Despite the additional insights these analyses revealed, the fundamental issue of the extent to which the defined AHP factors adequately encompass the vulnerability space remained.
For ANP, the initial values of the “relative importance of subcriteria with respect to criteria (W32)” submatrix were collected from the Nested AHP analysis. To test whether the nonlinearity of this submatrix significantly affected the ANP results, we changed the matrix values from Nested AHP to original AHP. There were only minor changes from the initial ANP results (Supplemental Figure. S4), suggesting that the nonlinear results of ANP were independent of the submatrix values.
Fig 4. These figures show the Nested AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 5. These figures show the ANP’s vulnerability score for the scope of our analysis, the lower values are scored as the least vulnerable areas while the ones with higher scores were ranked as the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 6. This graph shows the normalized amount each factor class 1, 2, 3, 4, 5 affects the overall weighted overlay for AHP, Nested AHP and ANP, where blue (circles) represents AHP, red (squares) represents Nested AHP and green (triangles) represents ANP.
3.3 1-N AHP
When we ran 1-N AHP for the worst case, N = 0.33, we attributed more weightage to the unknown acute and chronic factors than for the average case, N = 0.165. However, due to the low probability of the presence of an acute or chronic factor, the vulnerability scores were lower for the worst case than the average case (Fig 7 and Fig 8). Even when we considered the distribution of the result values, we saw that the results for the worst case were more right skewed than the average case, which was more right skewed than the original AHP. From this we can gather that given our assumptions, there was no negative impact of considering unknown factors in our AHP analysis. However, 1-N would positively be able to capture the robustness of AHP results in other contexts, or if the assumptions were altered, and since our acute and chronic layers were purely random across the Ganga buffer, the change in vulnerability values were correlated to the values of the initial AHP results, and not to any specific remote sensing or other GIS data layers in the analysis.
Fig 7. These figures show the 1 - N AHP worst case’s vulnerability score when we consider the worst case (N=0.33); the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 8. These figures show the 1 - N AHP average case’s vulnerability score when we consider the average/expected case (N=0.165); the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
3.4 Fuzzy AHP and Fuzzy 1-N AHP
Out of 100,000 Fuzzy AHP simulations, there were 4,433 rank reversals and no simulation had any second-order rank reversal (cases with greater than 1 index being flipped e.g. Population Density and Rainfall) at 95% fuzziness. This implied that there was a 4.43% chance of rank reversals at 95% fuzziness. Of these reversals, the vast majority occurred between the factors Rainfall and Drainage Density (Table 5), which can be attributed to the low percentage difference between their initial values.
Table 5: Count of Rank Reversals for Fuzzy AHP out of 100,000 simulations
Reversals between Parametersa Range of Original AHP results Count of Rank Reversals
PD and LULC (0.408, 0.268) 45
LULC and Rainfall (0.268, 0.138) 30
Rainfall and DD (0.138, 0.1) 4233
DD and Slope (0.1, 0.058) 146
Slope and LST (0.058, 0.028) 49
aHere, PD is Population Density, LULC is Land Use Land Cover, DD is drainage density, and LST is Land Surface Temperature
The simulations with rank reversals were tested to determine their potential impact on vulnerability scores. This was done by running AHP on a sample of 500 from the 4,433 cases where a rank reversal occurred, and averaging all the output layers to return the Mean Fuzzy AHP layer (Fig 9). Of these, areas that experienced changes in vulnerability due to rank reversals had high Drainage Density, or were areas that tended to receive large amounts of precipitation.
Fig 9. These figures show the Mean Fuzzy AHP’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
For both the cases, we saw the distribution skewed to the left as compared to the original AHP (Fig 10 and Fig 11), and the vulnerability increased, which was more prominent for Case 2 (Fig 11). Due to the high weightage of the parameter, there was a increase in vulnerability scores in the areas with high population density.
Fig 10. These figures show the Fuzzy AHP Case 1 vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
Fig 11. These figures show the Fuzzy AHP Case 2’s vulnerability score for the scope of our analysis; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
For Fuzzy 1-N AHP, there were 15,413 rank reversals out of 100,000 AHP simulations, suggesting a 15.4% chance of rank reversals (Table 6). There were around 200 second-order rank reversals as well (Table 7). Most of the rank reversals were between the factors Slope and Chronic, followed by Rainfall and DD, similar to Fuzzy AHP. A major contributor to the rank reversals was the closeness of the original AHP weights, leading to a higher likelihood of rank reversals when fuzziness is incorporated.
Table 6: Count of First-order Rank Reversals for Fuzzy 1-N AHP out of 100,000 simulations
Reversals between Parameters Range of Original AHP results Count of Rank Reversals
Acute and PD (0.297, 0.27336) 185
PD and LULC (0.27336, 0.17956) 854
LULC and Rainfall (0.17956, 0.09246) 21
Rainfall and DD (0.09246, 0.067) 6054
DD and Slope (0.067, 0.03886) 966
Slope and Chronic (0.03886, 0.033) 7220
Chronic and LST (0.033, 0.01876) 447
Table 7: Count of Second-order Rank Reversals for Fuzzy 1-N AHP out of 100,000 simulations
Reversals between Parameters Range of Original AHP results Count of Rank Reversals
Acute and LULC (0.297, 0.17956) 0
PD and Rainfall (0.27336, 0.09246) 0
LULC and DD (0.17956, 0.067) 0
Rainfall and Slope (0.09246, 0.03886) 5
DD and Chronic (0.067, 0.033) 77
Slope and LST (0.03886, 0.01876) 117
3.5 Comparison of analysis scores
To compare different approaches and test areas where the predicted vulnerability changes, we normalized the different output layers by standardizing them to follow Normal distributions with a Mean of 0 and Standard Deviation of 1, i.e. Standard Normal Distribution. This enabled us to compare different approaches that had different ranges. Then we subtracted each output layer from the original AHP results, i.e. AHP Vulnerability Score - Variant Vulnerability Score, to get our final difference layers. While computing these difference layers, the positive values showed that as per AHP, the areas were identified as more vulnerable than the variant map did, and the vice versa in case of negative values.
We compared the results of every other approach with the original AHP results, using AHP as the baseline for checking the robustness of our approaches. We wanted to infer three things from the difference layers, 1) Which pixel values changed relative to AHP, 2) How much did they change, 3) Were the changed values higher or lower than AHP?
To visualize and compare the difference of the AHP layer with other approaches, we created 3 different map types (Fig 12). We created Map Type 1 by stretching the colors between the overall min max of all the different approaches, to compare the approaches among themselves. For Map Type 2 we stretched the colors between the local minimum and maximum value of each approach, to make the outliers for each individual map stand out. And for Map Type 3, we stretched the colors between the local 5 percentile and 95 percentile for each approach, to prevent the outliers from washing-out the differences in the map.

### v3

@@ 5.2.29 Problem structuring N
The importance of factors influencing naala vulnerability were placed in the following order: Population Density (PD) > LULC > Rainfall (RAIN) > Drainage Density (DD) > Slope > LST, and the factor importance rankings were inputs into an AHP pairwise comparison matrix (S3 Table.).
S3 Table. contains the pairwise comparison matrix and the Normalized AHP pairwise comparison matrix. The Principal Eigenvalue is the mean of the eigenvectors, and n is the number of factors - six.
To begin the AHP process, we constructed a pairwise comparison matrix (A), where each element $a_{ij}$ represented the importance of criterion $i$ relative to criterion $j$.
@@
In some situations, AHP importance rankings may produce inconsistent outcomes due to matrix properties. The Consistency Index ($CI$) and Consistency Ratio ($CR$) were developed to test for inconsistency with a Consistency Ratio < 0.1 indicating the rankings were consistent [7]:
@@
In our case, with a consistency ratio ($CR$) of 0.06, the matrix was considered consistent (as $CR$ < 0.1). This consistency indicated that our pairwise comparison matrix was reliable for determining the weights.
S2 Text
Natural breaks and AHP factor classes Data for each AHP factor were divided into five vulnerability classes using the Jenks natural breaks method [19]. The Jenks method is widely used within GIS packages to generate variance-minimization classification. Breaks were typically uneven, and were selected to separate values where large changes in value occur. For this, we used the boxplots (S1 Fig.) and histograms (S2 Fig.) for the layers to identify the natural breaks for each factor.
S1 Fig.
AHP Box and Whisker Plot Summary Panels (a) through (e) are box and whisker plots for the AHP factors Population Density (a), Annual Rainfall (b), Drainage Density (c), Slope (d), and Land Surface Temperature (e) for the length of the Ganga contained in the area of analysis. Data distribution was tested and results are given for each factor label. Whiskers show the range of values, the box "shoulders" are the inter-quartile range, and the median is shown as a band.
S2 Fig.
AHP Histogram Summary Panels (a) through (l) are the histograms and log-histograms to assess the distributions of the AHP factors along the length of the Ganga contained in the area of analysis. Panels (a), (c), (e), (g), (i) represent the histograms for Population Density, Annual Rainfall, Drainage Density, Slope, and Land Surface Temperature, respectively, while panels (b), (d), (f), (h), (j) respectively represent the log-histograms for the same. Panels (a) through (j) are binned using deciles, and coloured green to red for low to high vulnerability. (k) and (l) represent LULC and are color coded to their respective Land Use classes.
S3 Text
Nested AHP pairwise comparisons and calculation
@@
Using a pairwise comparison matrix, we assessed the relationship between individual parameter values. S4 Table. shows the pairwise comparison of Population Density and the weight each category has, e.g. PD of >4036 people per km2 has a 9 times greater effect on vulnerability than when PD was <865 people per km2. The pairwise matrices for the other factors alongside the explanations of the comparison indices is as follows:
Minimum and maximum values for the five factor classes of population density, slope, rainfall, and drainage density were averaged, normalized, transformed to a 9-point scale to create the factor rankings, and fit to the Saaty AHP scale based on their relationships with pollution production or its transport as well as their breadth across a 9-point scale ( S3 Text and S4 Table.). The relationship between population density and pollution production was assumed linear based on municipal solid waste production in India [42], and the factors were ranked as 1, 4, 6, 8, and 9. Rainfall factor rankings spanned 1 through 5 on the Saaty scale because the range between lowest and highest average annual precipitation was a factor of 2 (1, 3, 4, 5, 5). Drainage density can have a nonlinear relationship with pollution transport [12], and although the range of drainage density values was more than 20-fold, its factors were ranked as 1, 6, 8, 9, and 9. The relationship between slope and transport is nonlinear for slopes greater than  26 degrees, and linear for slopes less than 26 degrees [43]. So slope factor rankings were ranked 1, 1.5, 3, 6, and 9. Minimum and maximum temperature data per factor rank were used to calculate normalized temperature dependent microbial growth rates with a Q10 (temperature coefficient) of 1.75. Then the growth rates were transformed to a 9-point scale and scored on the Saaty scale (1, 4, 5, 6, 8). LULC factor ranks were based on ESA WoldCover map [38] categories (built-up, cropland, bare/sparsely vegetated, tree cover/grassland, water/wetlands) (1,3,4,5,7). Built-up was treated as the factor rank 1 because of the breadth of pollution sources (e.g., municipal solid wastes, pathogens, excess nutrients, and heavy metals) and amount of impervious surfaces that lead to pollution transport. Cropland was ranked 3 due to runoff of fertilizers and agrochemicals. Bare/sparsely vegetated land was ranked 4 due to moderate transport risk. Tree cover and grassland were ranked 5 on the Saaty scale because these categories represent natural ecosystems, and water/wetlands were ranked as 7 on the Saaty scale, because Ganga was the point of pollution deposition.
@@
Running AHP on this pairwise matrix for Population Density gave the subweights for classes 1 - 5. Similarly, once we ran this for the other factors; LULC, Slope, Precipitation, Drainage Density, and Temperature, we got the priority vector of subweights for each factor (S4 Table.).
@@
  • $S$ is the 5x6 matrix of subweights, where each column corresponds to a factor and each row corresponds to the subweights for factor classes 1, 2, 3, 4, 5.
@@
S4 Table.
Nested AHP: Factor Vulnerability Rankings for all Factors LULC Built-up Cropland Barren/ Sparse Vegetation Trees/ Grass/ Shrubs Water/ Wetland Built-up 1 3 4 5 7 Cropland 0.33 1 1 2 4 Barren/Sparse Vegetation 0.25 1 1 1 3 Trees/ Grass/ Shrubs 0.2 0.5 1 1 2 Water / Wetland 0.14 0.25 0.33 0.5 1 Total 1.93 5.75 7.33 10 17 Rainfall (mm/y) >1410 1189-1410 1019-1189 868-1019 <868 >1410 1 3 4 5 5 1189-1410 0.33 1 1 2 2 1019-1189 0.25 1 1 1 1 868-1019 0.2 0.5 1 1 1 <868 0.2 0.5 1 1 1 Total 1.98 6 8 10 10 Drainage density (km/km2) >4 2.6-4 1.5-2.6 0.6-1.5 <0.6 >4 1 6 8 9 9 2.6-4 0.17 1 2 3 3 1.5-2.6 0.13 0.5 1 1 1 0.6-1.5 0.11 0.33 1 1 1 <0.6 0.11 0.33 1 1 1 Total 1.51 8.17 13 15 15 Slope (Angle %) >35 24-35 45650 45394 <4 >35 1 1.5 3 6 9 24-35 0.67 1 2 5 8 45650 0.33 0.67 1 4 7 45394 0.17 0.22 0.25 1 3 <4 0.11 0.13 0.14 0.33 1 Total 2.28 3.52 5.89 16 28 Temperature (℃) >37 34-37 30-34 25-30 <25 >37 1 4 5 6 8 34-37 0.25 1 1 2 4 30-34 0.2 1 1 1 3 25-30 0.17 0.5 1 1 2 <25 0.13 0.25 0.33 0.5 1 Total 1.74 6.75 8.33 11 18
S4 Text
Fuzzy AHP calculation
@@
In Fuzzy AHP, the decision maker generally recreates the pairwise matrix by choosing the $l,m,u$ for each element, but to understand the outlying cases for Fuzzy AHP, and to incorporate the breadth of decision maker’s range, we ran Fuzzy AHP on 100,000 Monte Carlo simulations at 95% fuzziness. For each run, we made a unique Fuzzy Pairwise Comparison Matrix, where if the comparison index between two factors was 2, the triangular fuzzy range would be (0.052,2,21.95)=(0.1,2,3.9)(0.05*2,2,2*1.95)=(0.1,2,3.9), and we chose a random value between 0.1 and 2, and 2 and 3.9 to design the final TFN [random.uniform(0.1,2),2,random.uniform(2,3.9)][random.uniform(0.1,2),2,random.uniform(2,3.9)]. Running 100,000 simulations took into account most of the likely Fuzzy Pairwise Comparison Matrices that might be generated.
@@
S5 Text
ANP supermatrix and inner dependence (W22 and W33) pairwise comparisons Pairwise comparisons were made to describe the relative importance of the ANP criteria (AHP factors) on influencing each other (S3 Fig.). Criteria that did not contribute to an inner dependence matrix were scored with a zero. When identifying the relative importance of Land Use Land Cover (LULC) compared to other criteria on influencing Population Density (PD), it was scored with a 2 because of the importance of built-up and agriculture categories in LULC and their direct relationships with population densities. Rainfall and slope were scored with a 5 because historically there were natural limits to both that influenced urbanization and agriculture although this is changing [44, 45, 46, 47, 48]. Drainage Density (DD) was scored an 8 because the natural drainage density is modified by people to meet agricultural and societal requirements suggesting it has a minor influence on PD. When identifying the relative importance of PD compared to other criteria on influencing LULC, it was scored with a 2 as a reciprocal of the LULC-PD inner dependence. Rainfall and temperature are major drivers of land cover type, and as such each was scored with a 4. Slope and DD shape landscapes within land cover types, and each was scored with a 7. When identifying the relative importance of LULC compared to other criteria on influencing rainfall, it was scored with a 7 and PD was scored with an 8 because of the potential for built-up land to influence local precipitation [49]. Slope may influence the speed of runoff during rainfall and was scored with a 5. When identifying the relative importance of criteria on influencing DD, PD was scored a 2, LULC was scored a 4, and slope was scored a 7, because DD is the product of both anthropogenic and natural influences. When identifying the relative importance of criteria on influencing temperature, PD was scored a 5, and LULC was scored an 8 to account for potential urban heat island effects.
S3 Fig.
ANP Parameter Inner Dependencies Inner dependencies between ANP criteria (AHP factors) are shown as ovals. Arrows about the ovals indicate influence by a criterion on other criteria to its right and arrows below the ovals indicate influence by a criterion on other criteria to its left.
S6 Text
1-N AHP calculation
@@
We first took our original output of AHP, i.e., the weights of the known factors. Here, we assumed that our known factors would at least represent 67% of the importance while classifying the vulnerability variable. This is a fair assumption since through our preliminary analysis we can conclude that our factors were robust enough to represent at least 67% of the variability of the vulnerability variable.
@@
To evaluate the unknown variability captured between 1-N (where N lies between 0.67 and 1), we split it between acute and chronic. We assumed that since acute factors have a significant impact on vulnerability, 90% of the 1-N factor would be attributed to the impact of acute factors to vulnerability, while the leftover 10% would be attributed to the impact of chronic factors to vulnerability. For instance, in the worst case scenario if 1-N is 0.33 (i.e., the weightage of the unknown factors towards vulnerability is 33%), the 0.33 will be split into 0.297 for acute and 0.033 for chronic. The original six factors, with their weights 0.408, 0.268, 0.138, 0.100, 0.058, 0.028 will be scaled down based on N. If N is 0.67, we multiply the old weights by 0.67 so that they sum up to n, and we can accommodate the new acute and chronic weights. Here, 0.408*0.67 gives 0.27336, which would be lower than 0.297 for acute factors, which means the acute factors would have a higher impact on vulnerability but with a lower probability.
@@
S4 Fig.
@@
ANP with original AHP submatrix These figures show the ANP’s vulnerability score with the original AHP submatrix; the lower values pertained to the least vulnerable areas while the ones with higher scores were the most vulnerable. Map of exact values per pixel location (a) and histogram of binned ranges of values (b).
S5 Fig.